I can't figure out how to set up this equation the find all three sides. The Perimeter of a triangle is 7ft. The longest side of a triangle measures 20 in more than the shortest side. The middle side is 2 times the measure of the shortest side. Find the lengths of the three sides in inches.
Answer by mananth(16946)

let shortest side be x in
longest side = x+20in ==> x+20
middle side = 2x
x+2x+x+20=84
4x+20=84
4x=64
x=16
2x=32
x+20=36
16,32,36
are the sides